# Unveiling the Unexplored Decay Mode of a Light Charged Higgs Boson to an Off-Shell Top Quark and a Bottom Quark

### v1

@@ Abstract
The charged Higgs boson ($H^{\pm}$) with a mass below the top quark mass remains a viable possibility within the type-I two-Higgs-doublet model under current constraints. While previous LHC searches have primarily focused on the $H^{\pm}\to\tau\nu$ decay mode, the decay channel into an off-shell top quark and a bottom quark, $H^{\pm}\rightarrow t^{*}b$, is leading or subleading for $H^{\pm}$ masses between 130 and 170 GeV. This study investigates the discovery potential of future colliders for this off-shell decay mode through pair-produced charged Higgs bosons decaying via $H^{+}H^{-}\rightarrow t^{*}b\tau\nu\rightarrow bbjj\tau\nu$. We perform signal-to-background analyses at the HL-LHC and a prospective 100 TeV proton-proton collider, employing cut-flow strategies and the Boosted Decision Tree method. However, due to the softness of the $b$ jets, signal significances fall below detection thresholds at these facilities. Extending our study to a multi-TeV muon collider (MuC), we demonstrate that a 3 TeV MuC achieves high signal significance, surpassing the $5\sigma$ threshold with an integrated luminosity of 1 ab-1, assuming a 10% background uncertainty. Specifically, for $M_{H^{\pm}}=130$, 150, and 170 GeV, the significances are 13.7, 13.5, and 6.06, respectively. In contrast, a 10 TeV MuC requires 10 ab-1 to achieve similar results. Our findings highlight the critical role of the MuC in probing the new signal channel $H^{\pm}\rightarrow t^{*}b$, offering a promising avenue for future charged Higgs boson searches involving off-shell top quarks.
Higgs Physics, Beyond the Standard Model, Muon Collider
I Introduction
@@
One of the most promising BSM signals at high-energy colliders is the presence of a light charged Higgs boson with a mass below the top quark mass $m_{t}$. This possibility remains viable under current constraints within the type-I and type-X [^1] In type-II and type-Y 2HDM, the charged Higgs boson is tightly constrained to be as heavy as $M_{H^{\pm}}\gtrsim 800{\;{\rm GeV}}$ due to the measurements of the inclusive $B$-meson decay into $s\gamma$ Misiak:2020vlo . two-Higgs-doublet model (2HDM) Aoki:2009ha ; Branco:2011iw ; Craig:2013hca ; Wang:2022yhm ; Shen:2022yuo ; Kanemura:2022ldq ; Lee:2022gyf , three-Higgs doublet model Akeroyd:2018axd , next-to-2HDM Abouabid:2021yvw , lepton-specific Inert doublet model Han:2021gfu , and scalar-triplet model Ferreira:2021bdj .
@@
Significant efforts have been made to search for the light charged Higgs boson at the LHC and future lepton colliders. For the decay $H^{\pm}\rightarrow\tau^{\pm}\nu$, various production channels have been explored, such as $t\rightarrow H^{\pm}b$ Abbaspour:2018ysj ; Demir:2018iqo ; ATLAS:2018gfm ; Sanyal:2019xcp ; Sirunyan:2019hkq ; Ghosh:2022wbe , $pp\to H^{\pm}\varphi^{0}$ Kim:2023lxc , $pp\rightarrow H^{\pm}A$ Kanemura:2011kx , $pp\rightarrow H^{+}H^{-}$ Duarte:2024zeh , $cs/cb\rightarrow H^{\pm}$ Hernandez-Sanchez:2012vxa ; Hernandez-Sanchez:2020vax , and $W^{\pm*}W^{\pm*}\rightarrow H^{\pm}H^{\pm}$ Aiko:2019mww . Here, $\varphi^{0}$ denotes a new CP-even neutral Higgs boson. For the $H^{\pm}\rightarrow cb/cs$ mode, the production channel of $t\rightarrow H^{\pm}b$ has been considered ATLAS:2013uxj ; CMS:2015yvc ; Akeroyd:2018axd ; CMS:2018dzl ; ATLAS:2021zyv ; CMS:2020osd ; Akeroyd:2022ouy . The $H^{\pm}\rightarrow W^{\pm}\varphi^{0}/W^{\pm}A$ modes have been extensively studied for production channels such as $t\rightarrow H^{\pm}b$ Dermisek:2012cn ; Arhrib:2020tqk ; Hu:2022gwd ; Fu:2023sng , $pp\rightarrow H^{\pm}\varphi^{0}$ Arhrib:2017wmo ; Mondal:2021bxa ; Arhrib:2022inj ; Kim:2022hvh ; Kim:2022nmm ; Bhatia:2022ugu ; Li:2023btx ; Mondal:2023wib , $pp\rightarrow H^{\pm}A$ Arhrib:2021xmc ; Arhrib:2021yqf , $e^{+}e^{-}\rightarrow H^{+}H^{-}$ Kausar:2020ims , $pp\rightarrow H^{+}H^{-}$ Arhrib:2021xmc ; Arhrib:2021yqf ; Shen:2022yuo , $W^{\pm*}W^{\pm*}\rightarrow H^{\pm}H^{\pm}$ Arhrib:2019ywg , $pp\rightarrow H^{\pm}W^{\mp}$ Krab:2022lih , and $pp\rightarrow H^{\pm}hh$ Kang:2022mdy .
However, one important decay mode has been largely overlooked: the decay of a light charged Higgs boson into an off-shell top quark (denoted as $t^{*}$) and a bottom quark. In the type-I 2HDM, where the Yukawa couplings of $H^{\pm}$ are inversely proportional to $\tan\beta$, the decay $H^{\pm}\rightarrow t^{*}b$ becomes the leading mode for $135{\;{\rm GeV}}\lesssim M_{H^{\pm}}\lesssim m_{t}$, with $H^{\pm}\rightarrow\tau\nu$ as the second leading mode. Thus, it is of great significance to investigate the discovery potential of future high-energy colliders for this decay mode. For the production of charged Higgs bosons, we consider pair production, yielding the final state of $H^{+}H^{-}\rightarrow t^{*}b\tau\nu\rightarrow bbjj\tau\nu$. This approach is necessitated by the stringent constraints from searches for $t\rightarrow bH^{\pm}(\rightarrow\tau\nu)$, which limit the branching ratio $\text{Br}(t\rightarrow bH^{\pm})$ to below $\mathcal{O}(10^{-4})$ in the type-I 2HDM.
We will rigorously investigate the discovery potential of the HL-LHC and a 100 TeV $pp$ collider for the signal $H^{+}H^{-}\rightarrow t^{*}b\tau\nu\rightarrow bbjj\tau\nu$, employing both cut-flow strategies and the Boosted Decision Tree (BDT) method. The analysis reveals that these efforts are unfortunately unsuccessful due to the softness of the $b$ jets. Given these challenges, the focus of our research then shifts to the multi-TeV Muon Collider (MuC). This facility promises to offer a higher boost for the $b$ jets, potentially enhancing their detectability and opening new avenues for exploration.
@@
The MuC stands out as a powerful tool for BSM searches Capdevilla:2020qel ; Bandyopadhyay:2021pld ; Sen:2021fha ; Asadi:2021gah ; Huang:2021nkl ; Choi:1999kn ; Han:2020uak ; Han:2021udl ; Jueid:2021avn ; Han:2022edd ; Black:2022qlg ; Belfkir:2023vpo ; Jueid:2023zxx , thanks to its clean collision environment, higher energy reach, reduced beamstrahlung, and efficient energy use. The prospects of the MuC program have been significantly enhanced by recent advancements in addressing critical challenges, such as cooling muon beams Antonelli:2013mmk ; Antonelli:2015nla and reducing beam-induced backgrounds (BIB) Collamati:2021sbv ; Ally:2022rgk .
@@
Our study will conduct a signal-to-background analysis for two collider configurations: $\sqrt{s}=3\text{ TeV}$ with a total integrated luminosity of 1 $1\text{ ab}^{-1}$ and $\sqrt{s}=10\text{ TeV}$ with a total integrated luminosity of 10 $10\text{ ab}^{-1}$. The analysis aims to demonstrate that the entire mass range of $M_{H^{\pm}}\in[130,170]\text{ GeV}$ can achieve a high signal significance, surpassing the $5\sigma$ discovery threshold. These findings represent our main contributions to the study of the light charged Higgs boson.
@@
The paper is organized as follows. In Sec. II, we briefly review the type-I 2HDM with CP invariance and softly broken $Z_{2}$ parity. Based on the results of random scans incorporating theoretical and experimental constraints, we investigate the characteristic features of the allowed parameters and suggest the golden channel to probe the unexplored $H^{\pm}\rightarrow t^{*}b$ decay mode, $H^{+}H^{-}\rightarrow t^{*}b\tau\nu\rightarrow bbjj\tau\nu$. Section III deals with the signal-to-background analysis at the HL-LHC and a 100 TeV $pp$ collider, incorporating comprehensive cut-based analysis and the BDT. In Sec. IV, we turn to the multi-TeV MuC and perform the signal-to-background analysis. Conclusions are presented in Sec. V.
II Brief review of the light charged Higgs boson in type-I 2HDM
@@
$+\frac{1}{2}\lambda_{1}(\Phi^{\dagger}_{1}\Phi_{1})^{2}+\frac{1}{2}\lambda_{2}(\Phi^{\dagger}_{2}\Phi_{2})^{2}+\lambda_{3}(\Phi^{\dagger}_{1}\Phi_{1})(\Phi^{\dagger}_{2}\Phi_{2})+\lambda_{4}(\Phi^{\dagger}_{1}\Phi_{2})(\Phi^{\dagger}_{2}\Phi_{1})$
@@
In 2HDM, there are five distinct physical Higgs bosons: the lighter CP-even scalar $h$, the heavier CP-even scalar $H$, the CP-odd pseudoscalar $A$, and a pair of charged Higgs bosons $H^{\pm}$. The relationships between these Higgs states and the weak eigenstates described in Equation 3 are determined by two mixing angles, $\alpha$ and $\beta$, which can be found in Ref. Song:2019aav . The SM Higgs boson ${h_{\rm SM}}$ is a linear combination of $h$ and $H$, specifically as ${h_{\rm SM}}=\sin(\beta-\alpha)h+\cos(\beta-\alpha)H$.
According to the assignment of $Z_{2}$ parity to the right-handed fermions, the model has four variants, type-I, type-II, type-X, and type-Y. The mass of $H^{\pm}$ in type-II and type-Y is heavily constrained by the measurements of the inclusive $B$-meson decay into $s\gamma$, requiring $M_{H^{\pm}}\gtrsim 800{\;{\rm GeV}}$ Misiak:2020vlo . Only type-I and type-X allow for the existence of charged Higgs bosons lighter than the top quark. Therefore, we focus on type-I in this study.
We employ two popular conditions: the Higgs alignment limit for the SM-like Higgs boson Carena:2013ooa ; Celis:2013rcs ; Cheung:2013rva ; Bernon:2015qea ; Chang:2015goa ; Das:2015mwa ; Kanemura:2021dez and the mass degeneracy of $H$ and $A$ for the electroweak precision data Kanemura:2011sj ; Chang:2015goa ; Chen:2019pkq . The Higgs alignment limit precludes the decay channel $H^{\pm}\to W^{\pm}h$. Furthermore, we restrict our analysis to scenarios where the charged Higgs boson is lighter than $H$ and $A$. Under this condition, the charged Higgs boson decays exclusively into fermion pairs, ensuring that the $H^{\pm}\rightarrow t^{*}b$ mode maintains a significant branching ratio. Conversely, if $M_{H/A}<M_{H^{\pm}}$, the decay modes $H^{\pm}\rightarrow W^{\pm(*)}H/A$ become accessible [^2] For a detailed phenomenological study of the scenario where $M_{H/A}<M_{H^{\pm}}$, see Ref. Sanyal:2023pfs ., suppressing the branching ratio of $H^{\pm}\rightarrow t^{*}b$.
@@
$\mathscr{L}_{\rm Yuk}=-\frac{1}{\tan\beta}\left\{\dfrac{\sqrt{2}V_{ud}}{v}H^{+}\overline{u}\left(m_{u}{P}_{-}-m_{d}{P}_{+}\right)d-\dfrac{\sqrt{2}m_{\tau}}{v}H^{+}\overline{\nu}_{L}\tau_{R}+{\rm H.c.}\right\},$ (6)
where $P_{\pm}=(1\pm\gamma^{5})/2$. Because these Yukawa couplings have a common factor of $1/\tan\beta$, the branching ratio of $H^{\pm}\rightarrow f\bar{f}$ is independent of $\tan\beta$.
$\mathscr{L}_{\rm gauge}=$ $\leavevmode\nobreak\ i\left[eA_{\mu}+\frac{g(s_{W}^{2}-c_{W}^{2})}{2c_{W}}Z_{\mu}\right]\left(H^{+}\partial^{\mu}H^{-}-H^{-}\partial^{\mu}H^{+}\right)$ (7)
$+\left[\frac{g^{2}}{2}W^{-\mu}W^{+}_{\mu}+e^{2}A^{\mu}A_{\mu}+\frac{g^{2}(s_{W}^{2}-c_{W}^{2})^{2}}{4c_{W}^{2}}Z^{\mu}Z_{\mu}+\frac{eg}{c_{W}(s_{W}^{2}-c_{W}^{2})}A^{\mu}Z_{\mu}\right]H^{+}H^{-},$
where $s_{W}=\sin\theta_{W}$, $c_{W}=\cos\theta_{W}$, and $\theta_{W}$ is the electroweak mixing angle.
To study the characteristics of the permissible parameter space for the light charged Higgs boson that predominantly decays into $t^{*}b$, we set $M_{H^{\pm}}=150{\;{\rm GeV}}$ and perform a random scan within the following parameter ranges:
$M_{H/A}\in[150,1000]{\;{\rm GeV}},\quad m_{12}^{2}\in[0,10^{5}]{\;{\rm GeV}}^{2},\quad\tan\beta\in[1,50].$ (8)

### v2

# Discovery Prospects for the Light Charged Higgs Boson Decay to an Off-Shell Top Quark and a Bottom Quark at Future High-Energy Colliders

@@ Abstract
The charged Higgs boson ($H^{\pm}$) with a mass below the top quark mass remains a viable possibility within the Type-I two-Higgs-doublet model under current constraints. While previous LHC searches have primarily focused on the $H^{\pm}\to\tau^{\pm}\nu$ decay mode, the decay channel into an off-shell top quark and a bottom quark, $H^{\pm}\rightarrow t^{*}b$, is leading or subleading for $H^{\pm}$ masses between 130 and 170 GeV. This study investigates the discovery potential of future colliders for this off-shell decay mode through pair-produced charged Higgs bosons decaying via $H^{+}H^{-}\rightarrow t^{*}b\tau\nu\rightarrow bbjj\tau\nu$. We perform signal-to-background analyses at the HL-LHC and a prospective 100 TeV proton-proton collider, employing cut-flow strategies and the Boosted Decision Tree method. However, due to the softness of the $b$ jets, signal significances fall below detection thresholds at these facilities. Extending our study to a multi-TeV muon collider (MuC), we demonstrate that a 3 TeV MuC achieves high signal significance, surpassing the $5\sigma$ threshold with an integrated luminosity of 1 ab-1 and a 10% background uncertainty. Specifically, for $M_{H^{\pm}}=130$, 150, and 170 GeV, the significances are 13.7, 13.5, and 6.06, respectively. In contrast, a 10 TeV MuC requires 10 ab-1 to achieve similar results. Our findings highlight the critical role of the MuC in probing the new signal channel $H^{\pm}\rightarrow t^{*}b$, offering a promising avenue for future charged Higgs boson searches involving off-shell top quarks.
pacs:
14.80.FdCharged Higgs bosons and 14.65.Hatop quarks
offprints:
1 Introduction
@@
One of the most promising BSM signals at high-energy colliders is the presence of a light charged Higgs boson with a mass below the top quark mass $m_{t}$. This possibility remains viable under current constraints within the Type-I and Type-X [^1] In Type-II and Type-Y 2HDM, the charged Higgs boson is tightly constrained to be as heavy as $M_{H^{\pm}}\gtrsim 800{\;{\rm GeV}}$ due to the measurements of the inclusive $B$-meson decay into $X_{s}\gamma$ Misiak:2020vlo . two-Higgs-doublet model (2HDM) Aoki:2009ha ; Branco:2011iw ; Craig:2013hca ; Wang:2022yhm ; Shen:2022yuo ; Kanemura:2022ldq ; Lee:2022gyf , three-Higgs doublet model Akeroyd:2018axd , next-to-2HDM Abouabid:2021yvw , lepton-specific Inert doublet model Han:2021gfu , and scalar-triplet model Ferreira:2021bdj .
A well-studied prototype of these BSM scenarios is the 2HDM, which extends the SM Higgs sector by introducing two complex $SU(2)_{L}$ Higgs doublets, $\Phi_{1}$ and $\Phi_{2}$, both carrying hypercharge $Y=+1$. These fields acquire vacuum expectation values, $v_{1}$ and $v_{2}$, whose ratio, $\tan\beta=v_{2}/v_{1}$, is a key parameter shaping the model’s phenomenology. The physical Higgs spectrum consists of five states: the lighter CP-even scalar $h$, the heavier CP-even scalar $H$, the CP-odd pseudoscalar $A$, and a pair of charged Higgs bosons, $H^{\pm}$. The existence of a light $H^{\pm}$ offers rich collider phenomenology, motivating dedicated experimental searches.
@@
Experimental efforts to discover a light charged Higgs boson span various colliders, including the LHC and future lepton colliders. For the decay $H^{\pm}\rightarrow\tau^{\pm}\nu$, various production channels have been explored, such as $t\rightarrow H^{+}b$ Abbaspour:2018ysj ; Demir:2018iqo ; ATLAS:2018gfm ; Sanyal:2019xcp ; Sirunyan:2019hkq ; Ghosh:2022wbe , $pp\to H^{\pm}\varphi^{0}$ Kim:2023lxc , $pp\rightarrow H^{\pm}A$ Kanemura:2011kx , $pp\rightarrow H^{+}H^{-}$ Duarte:2024zeh , $cs/cb\rightarrow H^{\pm}$ Hernandez-Sanchez:2012vxa ; Hernandez-Sanchez:2020vax , $pp\rightarrow W^{\pm*}W^{\pm*}\rightarrow H^{\pm}H^{\pm}$ Aiko:2019mww , and $e^{+}e^{-}\to H^{+}H^{-}$ Coniavitis:2007me . Here, $\varphi^{0}$ denotes a new CP-even neutral Higgs boson. For the $H^{\pm}\rightarrow cb/cs$ mode, the production channels of $t\rightarrow H^{+}b$ ATLAS:2013uxj ; CMS:2015yvc ; Akeroyd:2018axd ; CMS:2018dzl ; ATLAS:2021zyv ; CMS:2020osd ; Akeroyd:2022ouy and $e^{+}e^{-}\to H^{+}H^{-}$ Hou:2021qff have been considered. The decays $H^{\pm}\rightarrow W^{\pm}\varphi^{0}/W^{\pm}A$ have been extensively studied for production channels such as $t\rightarrow H^{\pm}b$ Dermisek:2012cn ; Arhrib:2020tqk ; Hu:2022gwd ; Fu:2023sng , $pp\rightarrow H^{\pm}\varphi^{0}$ Arhrib:2017wmo ; Mondal:2021bxa ; Arhrib:2022inj ; Kim:2022hvh ; Kim:2022nmm ; Bhatia:2022ugu ; Li:2023btx ; Mondal:2023wib ; Cheung:2022ndq , $pp\rightarrow H^{\pm}A$ Arhrib:2021xmc ; Arhrib:2021yqf , $e^{+}e^{-}\rightarrow H^{+}H^{-}$ Kausar:2020ims ; Barik:2022dvi ; Ouazghour:2023plc , $pp\rightarrow H^{+}H^{-}$ Arhrib:2021xmc ; Arhrib:2021yqf ; Shen:2022yuo , $pp\rightarrow W^{\pm*}W^{\pm*}\rightarrow H^{\pm}H^{\pm}$ Arhrib:2019ywg , $pp\rightarrow H^{\pm}W^{\mp}$ Krab:2022lih , $pp\rightarrow H^{\pm}hh$ Kang:2022mdy , and the associated productions of ${\ell^{+}\ell^{-}}\to H^{\pm}\tau^{\mp}\nu,tbH^{\pm},W^{\pm}H^{\mp}\varphi$ Ouazghour:2024twx .
Despite these extensive studies, the decay of a light charged Higgs boson into an off-shell top quark (denoted as $t^{*}$) and a bottom quark remains largely unexplored in terms of its discovery potential at future colliders. This decay mode is particularly significant in the Type-I 2HDM, where all Yukawa couplings of $H^{\pm}$ share a common $1/\tan\beta$ dependence. Consequently, the branching ratio for $H^{\pm}\rightarrow f\bar{f}^{\prime}$ is dictated primarily by fermion masses rather than $\tan\beta$ itself. If kinematically accessible, decays to heavier fermions dominate. In particular, $H^{\pm}\rightarrow t^{*}b$ emerges as the leading decay mode for $M_{H^{\pm}}\in[135{\;{\rm GeV}},m_{t}]$, with $H^{\pm}\rightarrow\tau^{\pm}\nu$ as the second most prominent channel. Even for $M_{H^{\pm}}$ slightly below 135 GeV, this mode remains among the most significant, highlighting its importance over the mass range $M_{H^{\pm}}\in[130,170]{\;{\rm GeV}}$. This motivates a detailed study of $H^{\pm}\to t^{*}b$ and its potential as a discovery channel at future colliders.
Previous research on $H^{\pm}\rightarrow t^{*}b$ has primarily focused on the branching ratios of the related $H^{\pm}\to b\bar{b}W^{\pm}$ decay, mainly within the Minimal Supersymmetric Standard Model, i.e., Type-II of the 2HDM Moretti:1994ds ; Djouadi:1995gv ; Borzumati:1998xr ; Bi:1999hp ; Djouadi:2005gj . Although Ref. Ma:1997up demonstrated that this decay mode could be enhanced for $M_{H^{\pm}}>140{\;{\rm GeV}}$ and $\tan\beta\lesssim 1$, their analysis was restricted to parton-level and focused only on the Tevatron.
To advance charged Higgs boson searches, it is essential to thoroughly investigate the observability of the $H^{\pm}\to t^{*}b$ decay mode at future high-energy colliders. This need forms the primary motivation for our work. We provide the first comprehensive analysis of the discovery potential for the $H^{\pm}\to t^{*}b$ decay mode across multiple future collider scenarios.
For the production of $H^{\pm}$ at high-energy colliders, we focus on pair production. This process depends solely on $M_{H^{\pm}}$, making it a robust and unambiguous choice for our study. Traditional searches use the channel $pp\to t\bar{t}$, followed by $t\to bH^{+}$ and $H^{\pm}\to\tau^{\pm}\nu$ Abbaspour:2018ysj ; Demir:2018iqo ; ATLAS:2018gfm ; Sanyal:2019xcp ; Sirunyan:2019hkq ; Ghosh:2022wbe , but this approach becomes highly inefficient in the large $\tan\beta$ limit. Current bounds already favor $\tan\beta$ values above about 6, which suppress $\text{Br}(t\to bH^{+})$ to below approximately $2\times 10^{-3}$ for $M_{H^{\pm}}=150{\;{\rm GeV}}$. If $\tan\beta=10$, this branching ratio further reduces to around $7\times 10^{-4}$. These values were obtained from our calculation using 2HDMC Eriksson:2009ws . Other production channels involving decays of additional BSM Higgs bosons are sensitive to their masses and $\tan\beta$. In contrast, pair production avoids these model-dependent complications, providing a clearer path for investigation.
@@
Motivated by these considerations, we will rigorously investigate the discovery potential of the HL-LHC and a 100 TeV $pp$ collider for the signal $H^{+}H^{-}\rightarrow t^{*}b\tau\nu\rightarrow bbjj\tau\nu$, employing both cut-flow strategies and the Boosted Decision Tree (BDT) method. However, the analysis reveals that these efforts are ultimately unsuccessful due to the softness of the $b$ jets. Given these challenges, our research shifts focus to the multi-TeV muon collider (MuC). The MuC is a powerful tool for BSM searches Capdevilla:2020qel ; Bandyopadhyay:2021pld ; Sen:2021fha ; Asadi:2021gah ; Huang:2021nkl ; Choi:1999kn ; Han:2020uak ; Han:2021udl ; Jueid:2021avn ; Han:2022edd ; Black:2022qlg ; Belfkir:2023vpo ; Jueid:2023zxx , offering a clean collision environment, higher energy reach, reduced beamstrahlung, and efficient energy usage. The prospects of the MuC program have been significantly enhanced by recent advancements in addressing critical challenges, such as cooling muon beams Antonelli:2013mmk ; Antonelli:2015nla and reducing beam-induced backgrounds (BIBs) Collamati:2021sbv ; Ally:2022rgk . The multi-TeV collision energy at the MuC promises a higher boost for the $b$ jets, potentially enhancing their detectability and opening new avenues for exploration.
@@
In our study, we conduct a signal-to-background analysis for two MuC configurations: $\sqrt{s}=3\text{ TeV}$ with a total integrated luminosity of 1 $1\text{ ab}^{-1}$ and $\sqrt{s}=10\text{ TeV}$ with a total integrated luminosity of 10 $10\text{ ab}^{-1}$. This analysis aims to demonstrate that the full mass range of $M_{H^{\pm}}\in[130,170]\text{ GeV}$ can achieve high signal significance, surpassing the $5\sigma$ discovery threshold. These findings represent our main contributions to the study of the light charged Higgs boson.
@@
The paper is organized as follows. In Sec. 2, we briefly review the Type-I 2HDM with CP invariance and softly broken $Z_{2}$ parity. Based on the results of random scans incorporating theoretical and experimental constraints, we investigate the characteristic features of the allowed parameters and suggest the golden channel to probe the unexplored $H^{\pm}\rightarrow t^{*}b$ decay mode, $H^{+}H^{-}\rightarrow t^{*}b\tau\nu\rightarrow bbjj\tau\nu$. Section 3 deals with the signal-to-background analysis at the HL-LHC and a 100 TeV $pp$ collider, incorporating comprehensive cut-based analysis and the BDT. In Sec. 4, we turn to the multi-TeV MuC and perform the signal-to-background analysis. Conclusions are presented in Sec. 5.
2 Brief review of the light charged Higgs boson in Type-I 2HDM
@@
$+\frac{1}{2}\lambda_{1}(\Phi^{\dagger}_{1}\Phi_{1})^{2}+\frac{1}{2}\lambda_{2}(\Phi^{\dagger}_{2}\Phi_{2})^{2}$
$+\lambda_{3}(\Phi^{\dagger}_{1}\Phi_{1})(\Phi^{\dagger}_{2}\Phi_{2})+\lambda_{4}(\Phi^{\dagger}_{1}\Phi_{2})(\Phi^{\dagger}_{2}\Phi_{1})$
@@
In 2HDM, there are five physical Higgs bosons: the lighter CP-even scalar $h$, the heavier CP-even scalar $H$, the CP-odd pseudoscalar $A$, and a pair of charged Higgs bosons $H^{\pm}$. These physical states are related to the weak eigenstates through the following transformations Song:2019aav :
$\left(\begin{array}[]{c}h_{1}\\ h_{2}\end{array}\right)$ $=$ $\mathbb{R}(\alpha)\left(\begin{array}[]{c}H\\ h\end{array}\right),\quad\left(\begin{array}[]{c}\eta_{1}\\ \eta_{2}\end{array}\right)=\mathbb{R}(\beta)\left(\begin{array}[]{c}z^{0}\\ A\end{array}\right),$ (13)
$\left(\begin{array}[]{c}w_{1}^{\pm}\\ w_{2}^{\pm}\end{array}\right)$ $=$ $\mathbb{R}(\beta)\left(\begin{array}[]{c}w^{\pm}\\ H^{\pm}\end{array}\right),$ (18)
where $z^{0}$ and $w^{\pm}$ are the Goldstone bosons that are absorbed to become the longitudinal components of the $Z$ and $W$ bosons, respectively. The rotation matrix $\mathbb{R}(\theta)$ is defined as
$\mathbb{R}(\theta)=\left(\begin{array}[]{cr}\cos\theta&-\sin\theta\\ \sin\theta&\cos\theta\end{array}\right).$ (21)
The SM Higgs boson ${h_{\rm SM}}$ is a linear combination of $h$ and $H$, specifically as ${h_{\rm SM}}=\sin(\beta-\alpha)h+\cos(\beta-\alpha)H$. In this study, we focus on the normal scenario where $h$ is the observed Higgs boson with a mass of 125 GeV.
According to the assignment of $Z_{2}$ parity to the right-handed fermions, the model has four variants, Type-I, Type-II, Type-X, and Type-Y. The mass of $H^{\pm}$ in Type-II and Type-Y is heavily constrained by the measurements of the inclusive $B$-meson decay into $X_{s}\gamma$, which requires $M_{H^{\pm}}\gtrsim 800{\;{\rm GeV}}$ Misiak:2020vlo . Only Type-I and Type-X allow for the existence of charged Higgs bosons lighter than the top quark. Since we aim to study the $H^{\pm}\to t^{*}b$ decay, we focus on the Type-I model.
The model has six physical parameters of $M_{H^{\pm}}$, $M_{H}$, $M_{A}$, $m_{12}^{2}$, $\tan\beta$, and $\sin(\beta-\alpha)$. To focus on the key features, we employ two popular conditions: the Higgs alignment limit for the SM-like Higgs boson Carena:2013ooa ; Celis:2013rcs ; Cheung:2013rva ; Bernon:2015qea ; Chang:2015goa ; Das:2015mwa ; Kanemura:2021dez and the mass degeneracy of $H$ and $A$ for the electroweak precision data Kanemura:2011sj ; Chang:2015goa ; Chen:2019pkq . The Higgs alignment limit precludes the decay channel $H^{\pm}\to W^{\pm}h$. Furthermore, we restrict our analysis to the scenario where the charged Higgs boson is lighter than $H$ and $A$ [^2] If $M_{H/A}<M_{H^{\pm}}$, the decay modes $H^{\pm}\rightarrow W^{\pm(*)}H/A$ become accessible Sanyal:2023pfs . In this scenario, the charged Higgs boson decays exclusively into fermion pairs, ensuring that the $H^{\pm}\rightarrow t^{*}b$ mode maintains a significant branching ratio.
@@
Type-I: $m_{h}=125{\;{\rm GeV}},\quad\sin(\beta-\alpha)=1,$
$M_{H^{\pm}}<M_{H/A}(=M_{H}=M_{A}).$
In this setup, the quartic couplings in section 2 are Das:2015mwa :
$\lambda_{1}$ $=$ $\frac{1}{v^{2}}\left[\left(\tan\beta\right)^{2}(M_{H}^{2}-M^{2})-m_{h}^{2}\right],$ (23)
$\lambda_{2}$ $=$ $\frac{1}{v^{2}}\left[m_{h}^{2}+\frac{1}{\tan\beta^{2}}\left(M_{H}^{2}-M^{2}\right)\right],$
$\lambda_{3}$ $=$ $\frac{1}{v^{2}}\left[m_{h}^{2}+2M_{H^{\pm}}^{2}-M_{H}^{2}-M^{2}\right],$
$\lambda_{4}$ $=$ $\frac{1}{v^{2}}\left[M_{A}^{2}-2M_{H^{\pm}}^{2}+M^{2}\right],$
$\lambda_{5}$ $=$ $\frac{1}{v^{2}}\left[M^{2}-M_{A}^{2}\right],$
where $M^{2}=m_{12}^{2}/(s_{\beta}c_{\beta})$.
$\mathscr{L}_{\rm Yuk}$ $=-\frac{1}{\tan\beta}\left\{\dfrac{\sqrt{2}V_{ud}}{v}H^{+}\overline{u}\left(m_{u}{P}_{-}-m_{d}{P}_{+}\right)d\right.$ (24)
$\left.-\dfrac{\sqrt{2}m_{\tau}}{v}H^{+}\overline{\nu}_{L}\tau_{R}+{\rm H.c.}\right\},$
where $P_{\pm}=(1\pm\gamma^{5})/2$. Given that these Yukawa couplings share a common factor of $1/\tan\beta$, the branching ratio of $H^{\pm}\rightarrow f\bar{f}$ depends solely on the fermion mass, independent of $\tan\beta$. As a result, when kinematically allowed, the decay of $H^{\pm}\to t^{(*)}b$ becomes overwhelmingly dominant.
@@
$+H^{+}H^{-}\left[\frac{g^{2}}{2}W^{-\mu}W^{+}_{\mu}+e^{2}A^{\mu}A_{\mu}\right.$
$\left.+\frac{g^{2}(s_{W}^{2}-c_{W}^{2})^{2}}{4c_{W}^{2}}Z^{\mu}Z_{\mu}+\frac{eg}{c_{W}(s_{W}^{2}-c_{W}^{2})}A^{\mu}Z_{\mu}\right],$
[…]
To study the characteristics of the permissible parameter space for the light charged Higgs boson that predominantly decays into $t^{*}b$, we consider three cases of $M_{H^{\pm}}=130,~{}150,~{}170{\;{\rm GeV}}$ and perform a random scan within the following parameter ranges:
$M_{H/A}$ $\in[150,1000]{\;{\rm GeV}},\quad m_{12}^{2}\in[0,10^{5}]{\;{\rm GeV}}^{2},$ (26)
$\tan\beta$ $\in[1,50].$